# Words Matter: Leveraging Individual Text Embeddings for Code Generation in CLIP Test-Time Adaptation

### v2

@@ 4.1 Online pseudolabels assignments
The renormalization vectors in our setting are computed trough a small number of matrix multiplications using the iterative Sinkhorn-Knopp algorithm, [20], where in each iteration $\mathbf{u}^{(t)}_{m}=\mathbf{u}_{m}/((\exp(\mathbf{T}_{m}^{\top}\mathbf{Z}/\lambda)\mathbf{v}^{t-1}_{m})$ and $\mathbf{v}^{(t)}_{m}=\mathbf{v}_{m}/((\exp(\mathbf{T}_{m}^{\top}\mathbf{Z}/\lambda)\mathbf{u}^{t}_{m})$, with $\mathbf{v}^{0}_{m}=\mathbf{1}$. Furthermore, $\mathbf{T}_{m}=[\mathbf{t}_{1m},...,\mathbf{t}_{Km}]$ denotes the matrix containing the class text prototypes obtained from the $m$-th template. Note that the computational burden introduced by OT is negligible, representing roughly $1\%$ of the total runtime for a batch, as discussed in Section 5.3.
4.2 Updating CLIP through textual codes
@@
Distilling knowledge from multiple text prototypes.  To optimize the objective in Eq. 3, we employ the soft label assignments, as they yield superior performance compared to their hard counterpart in other problems [5]. Now, to distill the knowledge of the more diverse, and richer representations derived from multiple text predictions per category to the model, the cross-entropy in Eq. 3 is optimized $M$ times, each one using a different template $m$, which can be defined for each image and template as $\ell(\mathbf{p}_{i},\mathbf{q}^{*}_{im})=-\mathbf{q}^{*}_{im}\log\mathbf{p}_{i}$, where $\mathbf{p}_{i}$ is obtained with the average text template as in Eq. 1. Thus, for a given test image $i$ in a batch, after each update produced by the $m$-th text template, the visual encoder produces novel visual embeddings $\mathbf{z}_{i}$, which are then used to generate new predictions $\mathbf{p}_{i}$ for $i$.
@@
Whole learning strategy.  The final proposed algorithm consists in updating the affine parameters in Layer Normalization, such as in [39, 31, 14], and learning a label assignment matrix $\mathbf{Q}_{m}$ per text template by solving the optimization problem presented in Eq. (7). This is done iteratively by alternating two steps across the batches and text templates: i) with the learnable parameters fixed, we compute the visual features $\mathbf{Z}\in^{d\times B_{s}}$ of the test batch samples, and find $\mathbf{Q}_{m}^{*}$ through Eq. (9) by iterating the updates on $\mathbf{u}_{m}$ and $\mathbf{v}_{m}$; and ii) given the current label assignments $\mathbf{Q}_{m}^{*}$, and the softmax predictions for the test samples obtained with each average class prototype $\mathbf{t}_{k}$, the set of learnable parameters is updated by minimizing Eq. (3) w.r.t. the layer norm parameters, where stochastic gradient descent is applied over the whole batch. This process is repeated $M$ times. Finally, the class predictions on the batch are inferred with the updated model. A view of the proposed method is detailed in Algorithm 1.
Thus, in a nutshell, we generate multiple pseudo-labels per sample, each derived from one text template. Note that pseudo-labeling is a common strategy in other fields, such as self-supervised [49] or semi-supervised [2, 46] learning. Thus, while $\mathbf{Q}$, or pseudo-labels in general, may contain errors, using these to refine the model is a standard practice, which typically yields consistent improvements. We refer the reader to Appendix G for a detailed discussion of the benefits of our approach compared to pseudo-labeling in other learning scenarios.
Interpretation.  Given a batch of $B_{s}$ visual embeddings $\mathbf{Z}=[\mathbf{z}_{1},...,\mathbf{z}_{B}]$, we aim at optimizing a function to map these embeddings into the class prototypes $\mathbf{T}=[\mathbf{t}_{1},...,\mathbf{t}_{K}]$. If we let $\mathbf{Q}$ denote the codes (or mapping), our goal is optimizing $\mathbf{Q}$ to maximize the similarity between the image features and the class text prototypes. While in fully-supervised, or semi-supervised regimes, labeled data is available to generate the class prototypes, in our Test-Time Adaptation scenario this becomes more challenging, as no label information about the test images is provided. Nevertheless, in the context of large vision-language models, the text can be leveraged as a form of supervision to create the prototypes $\mathbf{T}$, which represent a robust representation of each class. Furthermore, the entropic constraint term controls the smoothness of the mapping, promoting equally distributed predictions over all classes in each batch, which avoids degenerate solutions. Lastly, resorting to multiple templates individually, instead of averaging their embeddings, fully leverage the diverse information present in each text prompt.
Algorithm 1 CLIP-OT adaptation for one corruption.
1:input: test dataset $\mathcal{X}=\{\mathbf{x}_{n}\}$ (images), text templates $\mathcal{T}$ (class descriptions), visual and text encoders ($\boldsymbol{\theta}$, $\phi$). // Split $\mathcal{X}$ into $B$ batches of size $B_{s}$. // Compute $K\times M$ text prototypes ($\boldsymbol{\phi}(\mathcal{T}),\forall m,k)$.
2:$\mathbf{T}\in\mathbb{R}^{d\times K\times M}=\left[\mathbf{t}_{km}\right]_{k=1,\dots,K,m=1,\dots,M}$
3:for sampled minibatch $\{\mathbf{x}_{i}\}_{i=1}^{B_{s}}$ do // 1 - Adaptation
4:     for each template $m$ in a random permutation of $\{1,2,\dots,M\}$ do    // Step 1: Generation of pseudo-codes $Q^{*}_{m}$.
5:         $\mathbf{T}_{m}\in\mathbb{R}^{d\times K}$ // class text embeddings for $m$.
6:         $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{s}}]$ // visual features ($\boldsymbol{\theta}_{\text{LN}}^{(m-1)}$).
7:         Compute codes $\mathbf{Q}^{*}_{m}$ // Eq. (9).
8:         // Step 2: Distill knowledge to the model.
9:         $\mathbf{t}_{k}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{t}_{km}\quad(\mathbf{t}_{k}\in\mathbb{R}^{d})$ // average class text embeddings over all M templates.
10:         $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{s}}]$ // predict: $\mathbf{t}_{k},\forall k$, Eq. (1).
11:         Min. cross-entropy with $\mathbf{P}$ and $\mathbf{Q}^{*}_{m}$ // Eq. (3).
12:         $\theta_{\text{LN}}^{(m-1)}\rightarrow\theta_{\text{LN}}^{(m)}$ // Update layer norm (LN) of $\boldsymbol{\theta}$.
13:     end for // 2 - Inference (for all images in the batch).
14:     $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{s}}]$ // visual features (with $\boldsymbol{\theta}_{\text{LN}}^{(m)}$).
15:     $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{s}}]$ // predict with $\mathbf{t}_{k},\forall k$, Eq. (1).
16:end for
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours $\Delta$
CIFAR-10 88.74 91.69 $\pm$0.10 88.06 $\pm$0.06 90.04 $\pm$0.13 91.41 $\pm$0.17 91.05 $\pm$0.06 93.15 $\pm$0.10 +2.10 $\blacktriangle$
CIFAR-10.1 83.25 87.60 $\pm$0.45 81.80 $\pm$0.27 86.35 $\pm$0.27 87.78 $\pm$0.05 86.98 $\pm$0.31 88.37 $\pm$0.10 +1.39 $\blacktriangle$
CIFAR-10C Gaussian Noise 35.27 41.27 $\pm$0.27 33.90 $\pm$0.08 59.90 $\pm$0.36 61.89 $\pm$0.24 63.84 $\pm$0.24 64.85 $\pm$0.14 +1.01 $\blacktriangle$
Shot Noise 39.67 47.20 $\pm$0.23 38.20 $\pm$0.02 62.77 $\pm$0.07 63.52 $\pm$0.08 65.28 $\pm$0.21 67.34 $\pm$0.21 +2.06 $\blacktriangle$
Impulse Noise 42.61 48.58 $\pm$0.31 37.66 $\pm$0.20 56.02 $\pm$0.16 57.13 $\pm$0.02 58.64 $\pm$0.11 62.27 $\pm$0.43 +3.63 $\blacktriangle$
Defocus Blur 69.76 77.12 $\pm$0.16 67.83 $\pm$0.28 76.74 $\pm$0.05 78.86 $\pm$0.09 78.94 $\pm$0.12 82.09 $\pm$0.05 +3.15 $\blacktriangle$
Glass Blur 42.40 52.65 $\pm$0.30 38.81 $\pm$0.12 61.77 $\pm$0.16 62.88 $\pm$0.06 65.12 $\pm$0.07 68.07 $\pm$0.02 +2.95 $\blacktriangle$
Motion Blur 63.97 71.25 $\pm$0.09 63.39 $\pm$0.13 76.01 $\pm$0.19 76.85 $\pm$0.26 77.81 $\pm$0.14 81.30 $\pm$0.18 +3.49 $\blacktriangle$
Zoom Blur 69.83 76.20 $\pm$0.19 68.95 $\pm$0.16 77.40 $\pm$0.20 79.35 $\pm$0.04 79.32 $\pm$0.07 83.13 $\pm$0.24 +3.81 $\blacktriangle$
Snow 71.78 78.29 $\pm$0.20 70.16 $\pm$0.10 77.29 $\pm$0.16 79.44 $\pm$0.09 79.79 $\pm$0.06 83.71 $\pm$0.17 +3.92 $\blacktriangle$
Frost 72.86 79.84 $\pm$0.09 72.39 $\pm$0.22 79.20 $\pm$0.08 80.13 $\pm$0.10 80.54 $\pm$0.12 83.40 $\pm$0.20 +2.86 $\blacktriangle$
Fog 67.04 77.39 $\pm$0.01 64.31 $\pm$0.28 75.74 $\pm$0.14 77.68 $\pm$0.07 78.53 $\pm$0.22 82.56 $\pm$0.08 +4.03 $\blacktriangle$
Brightness 81.87 87.78 $\pm$0.03 81.30 $\pm$0.18 86.59 $\pm$0.16 87.10 $\pm$0.10 87.11 $\pm$0.11 89.90 $\pm$0.06 +2.79 $\blacktriangle$
Contrast 64.37 79.47 $\pm$0.11 62.26 $\pm$0.31 77.82 $\pm$0.14 80.04 $\pm$0.24 81.20 $\pm$0.22 84.86 $\pm$0.11 +3.66 $\blacktriangle$
Elastic Transform 60.83 70.00 $\pm$0.25 56.43 $\pm$0.27 70.20 $\pm$0.01 71.76 $\pm$0.10 72.66 $\pm$0.15 76.08 $\pm$0.30 +3.42 $\blacktriangle$
Pixelate 50.53 63.74 $\pm$0.18 42.80 $\pm$0.40 66.52 $\pm$0.13 69.28 $\pm$0.09 71.11 $\pm$0.13 76.25 $\pm$0.05 +5.14 $\blacktriangle$
JPEG Compression 55.48 62.64 $\pm$0.14 53.67 $\pm$0.25 63.51 $\pm$0.14 66.49 $\pm$0.14 67.36 $\pm$0.28 70.03 $\pm$0.37 +2.67 $\blacktriangle$
Average 59.22 67.56 56.80 71.17 72.83 73.82 77.06 +3.24 $\blacktriangle$
CIFAR-100 61.68 69.74 $\pm$0.16 63.78 $\pm$0.28 69.79 $\pm$0.04 70.38 $\pm$0.14 70.74 $\pm$0.20 71.68 $\pm$0.09 +0.94 $\blacktriangle$
CIFAR-100C Gaussian Noise 14.80 14.38 $\pm$0.14 14.03 $\pm$0.10 25.32 $\pm$0.14 31.28 $\pm$0.03 32.07 $\pm$0.23 33.43 $\pm$0.13 +1.36 $\blacktriangle$
Shot Noise 16.03 17.34 $\pm$0.27 15.25 $\pm$0.17 27.90 $\pm$0.05 33.44 $\pm$0.11 34.36 $\pm$0.11 35.60 $\pm$0.19 +1.24 $\blacktriangle$
Impulse Noise 13.85 10.03 $\pm$0.13 13.01 $\pm$0.13 25.62 $\pm$0.09 29.40 $\pm$0.11 30.33 $\pm$0.03 30.94 $\pm$0.41 +0.61 $\blacktriangle$
Defocus Blur 36.74 49.05 $\pm$0.07 37.60 $\pm$0.17 49.88 $\pm$0.23 52.32 $\pm$0.28 52.99 $\pm$0.16 53.87 $\pm$0.22 +0.88 $\blacktriangle$
Glass Blur 14.19 3.71 $\pm$0.07 16.41 $\pm$0.02 27.89 $\pm$0.03 31.20 $\pm$0.12 32.15 $\pm$0.30 35.26 $\pm$0.22 +3.11 $\blacktriangle$
Motion Blur 36.14 46.62 $\pm$0.27 37.52 $\pm$0.23 47.93 $\pm$0.14 49.72 $\pm$0.15 50.53 $\pm$0.12 52.77 $\pm$0.07 +2.24 $\blacktriangle$
Zoom Blur 40.24 51.84 $\pm$0.15 42.99 $\pm$0.11 52.70 $\pm$0.06 54.72 $\pm$0.04 55.30 $\pm$0.22 56.71 $\pm$0.17 +1.41 $\blacktriangle$
Snow 38.95 46.71 $\pm$0.21 42.35 $\pm$0.13 49.72 $\pm$0.01 51.79 $\pm$0.04 52.77 $\pm$0.15 54.30 $\pm$0.19 +1.53 $\blacktriangle$
Frost 40.56 44.90 $\pm$0.27 43.31 $\pm$0.14 49.63 $\pm$0.12 53.04 $\pm$0.08 53.79 $\pm$0.31 54.92 $\pm$0.07 +1.13 $\blacktriangle$
Fog 38.00 47.31 $\pm$0.04 38.81 $\pm$0.17 48.77 $\pm$0.04 50.78 $\pm$0.24 51.49 $\pm$0.21 53.57 $\pm$0.02 +2.08 $\blacktriangle$
Brightness 48.18 60.58 $\pm$0.18 50.23 $\pm$0.11 61.27 $\pm$0.08 62.65 $\pm$0.25 63.57 $\pm$0.21 64.43 $\pm$0.31 +0.86 $\blacktriangle$
Contrast 29.53 45.90 $\pm$0.11 28.09 $\pm$0.09 48.55 $\pm$0.24 51.34 $\pm$0.10 52.76 $\pm$0.27 55.01 $\pm$0.29 +2.25 $\blacktriangle$
Elastic Transform 26.33 33.09 $\pm$0.08 28.12 $\pm$0.15 37.45 $\pm$0.08 39.97 $\pm$0.06 40.90 $\pm$0.43 43.79 $\pm$0.27 +2.89 $\blacktriangle$
Pixelate 21.98 26.47 $\pm$0.09 20.43 $\pm$0.14 33.88 $\pm$0.14 39.59 $\pm$0.09 40.97 $\pm$0.16 44.51 $\pm$0.32 +3.54 $\blacktriangle$
JPEG Compression 25.91 29.89 $\pm$0.07 28.82 $\pm$0.09 36.07 $\pm$0.32 38.99 $\pm$0.16 39.59 $\pm$0.08 40.83 $\pm$0.14 +1.24 $\blacktriangle$
Average 29.43 35.19 30.46 41.51 44.68 45.57 47.33 +1.76 $\blacktriangle$
@@
Link to unsupervised representation learning.  The proposed distilling strategy takes inspiration from unsupervised representation learning [7, 28, 5], where comparing random crops of the same image has the potential to yield stronger image representations. In particular, these strategies leverage these novel “views” of the same image as an image augmentation strategy, whose integration improves the semantic quality of the learned image representations. Nevertheless, i) this strategy substantially increases the computational and memory burden [7, 28], ii) specific composition of data augmentation operations is crucial for learning good representations [7], and iii) low-quality crops can alter the model performance [5]. In contrast, exploiting multiple text embeddings to guide the model adaptation process serves as a similar strategy, while also alleviating these shortcomings. More concretely, the text embeddings are low dimensional vectors, which can be computed off-line as they are image agnostic, considerably reducing computation requirements. Furthermore, in the absence of supervision, we argue that diverse text descriptions associated with a given class convey more informative representations than individual crops or “views” of an image to learn more generalizable features. While images provide valuable visual information, their parts may lack the contextual richness necessary for effective representation learning. This distinction suggests that leveraging a variety of textual prompts can indeed serve as a more promising alternative to contrastive learning based on multiple image crops while also being more efficient.
@@ 5.1 Setting
Datasets.  All methods are benchmarked on standard TTA datasets known for their complexity and task diversity. These datasets include corrupted modifications from well-known datasets proposed in [15], such as CIFAR-10C, CIFAR-100C, or Tiny-ImageNet-C. Also, we employ its non-corrupted versions, i.e., CIFAR-10 and CIFAR-100 [21], CIFAR-10.1 [34], and Tiny-ImageNet [41]. Additionaly, we include Domainbed datasets (PACS [24], OfficeHome [38], VisDA-C [32]). More details can be found in Appendix B.
Baselines.  We resort to relevant CLIP-based TTA parameter-efficient adaptation approaches, i.e., modify a set of learnable parameters, including TENT [39], TPT [36], CLIPArTT [14], and WATT [31] (both versions WATT-S and WATT-P).
Architectures.  We employed the Vision Transformer architecture as the backbone for our experiments. The choice of ViT is driven by its superior performance on image classification tasks due to its attention mechanisms, which effectively capture long-range dependencies and spatial relationships within images. If the opposite is not stated, experiments are carried out using ViT-B/32, although we also provide further evaluations with larger ViT-backbones, e.g., ViT-B/16.
Implementation details.  Following the existing literature [31, 14], we used a batch size of 128 in our experiments and a learning rate of $10^{-4}$. The only additional hyper-parameters specific to our formulation are the entropic constraint weight, $\epsilon$ in Eq. 7, and the number of repetitions in the Sinkhorn algorithm. These are fixed to the same values across all experiments. $\epsilon$ is set to $\epsilon=0.7$, whereas the latter is fixed to $T=3$, which was enough for convergence, as in previous works [3]. Finally, we ran the adaptation experiments for each corruption/dataset three times under different random seeds and averaged the obtained accuracy accordingly.

### v3

@@ 4.2 Semantic alignment with multiple templates
The renormalization vectors in our setting are computed trough a small number of matrix multiplications using the iterative Sinkhorn-Knopp algorithm [knight2008sinkhorn], where in each iteration $\mathbf{u}^{(t)}_{m}=\mathbf{u}_{m}/((\exp(\mathbf{T}_{m}^{\top}\mathbf{Z}/\lambda)\mathbf{v}^{t-1}_{m})$ and $\mathbf{v}^{(t)}_{m}=\mathbf{v}_{m}/((\exp(\mathbf{T}_{m}^{\top}\mathbf{Z}/\lambda)\mathbf{u}^{t}_{m})$, with $\mathbf{v}^{0}_{m}=\mathbf{1}$.
2
For Generalization, we need consistency. To ensure generalization, model predictions must remain consistent across semantic views, preventing the visual encoder from overfitting to any single template, or “view”. In the proposed adaptation strategy, the model’s prediction $p(y=k|\mathbf{x}_{i})$ in Eq. 1 is always computed using a single classification head, defined by the averaged text class prototypes $\mathbf{T}$. This design enforces stability across templates, ensuring that the pseudo-supervision $q(y=k|\mathbf{x}_{i})$ is always applied against a consistent prediction function, preventing the encoder from overfitting to template‑specific fluctuations and promoting robust generalization.
@@
Coupling both together. First, to optimize the objective in Eq. 3, we employ the soft label assignments from (9), as they yield superior performance compared to their hard counterpart in other problems [caron2020unsupervised]. Now, to distill the knowledge of the more diverse, and richer representations derived from multiple text predictions, the cross-entropy in Eq. 3 is optimized by iterating through each of the $M$ views, which can be formally defined for each image $i$ and template $m$ as $\ell(\mathbf{p}_{i},\mathbf{q}^{*}_{im})=-\mathbf{q}^{*}_{im}\log\mathbf{p}_{i}$, where $\mathbf{p}_{i}$ is obtained with the average text template. Thus, for a given test image $i$ in a batch, after each update produced by the $m$-th text template, the visual encoder produces novel visual embeddings $\mathbf{z}_{i}$, which are then used to generate new predictions $\mathbf{p}_{i}$ for $i$.
4.3 Whole learning strategy
@@
The complete Semantic Anchor Transport (SAT) algorithm (Alg. 1) consists in learning a label assignment matrix $\mathbf{Q}_{m}$ per text template by solving the optimization problem presented in Eq. (7), and then updating the affine parameters of the visual encoder’s Layer Normalization layers, such as [wangtent, osowiechi2024watt, hakim2024clipartt]. This is done iteratively by alternating two steps across the batches and text templates: i) with the learnable parameters fixed, we compute the visual features $\mathbf{Z}\in\real^{d\times B_{s}}$ of the test batch samples, and find $\mathbf{Q}_{m}^{*}$ through Eq. (9) by iterating the updates on $\mathbf{u}_{m}$ and $\mathbf{v}_{m}$; and ii) given label assignments $\mathbf{Q}_{m}^{*}$, and the softmax predictions for the test samples obtained with each average class prototype $\mathbf{t}_{k}$, the set of learnable parameters is updated by minimizing Eq. (3) w.r.t. the layer norm parameters, where stochastic gradient descent is applied over the whole batch. This process is repeated $M$ times. Finally, the class predictions on the batch are inferred with the updated model.
Algorithm 1 Semantic Anchor Transport for one test batch.
1:Input: Test batch $\{\mathbf{x}_{i}\}_{i=1}^{B_{T}}$, set of $M$ semantic anchor matrices $\{\mathbf{T}_{m}\}_{m=1}^{M}$, visual encoder $\boldsymbol{\theta}$.
2:// $\mathbf{T}_{m}$ (and thus $\mathbf{T}$) are pre-computed offline from a systematically generated set of text templates.
@@
4:// — Adaptation Phase —
5:for each template $m$ in a random permutation of $\{1,\dots,M\}$ do
6:  // Step 1: Align - Compute soft assignments that maximize cross-modal alignment for the m-th “view”
7:  $\mathbf{Z}\leftarrow[\boldsymbol{\theta}(\mathbf{x}_{1}),\dots,\boldsymbol{\theta}(\mathbf{x}_{B_{T}})]$
8:  $\mathbf{Q}^{*}_{m}\leftarrow\text{SolveOT}(\mathbf{Z},\mathbf{T}_{m})\quad$ Eq. 9
9:  // Step 2: Adapt - Distill knowledge into a general representation
10:  $\mathbf{P}\leftarrow\text{Predict}(\mathbf{Z},\mathbf{T})\quad$ Eq. 1
11:  $\mathcal{L}\leftarrow\text{CrossEntropy}(\mathbf{P},\mathbf{Q}^{*}_{m})$
12:  Update LayerNorm parameters of $\boldsymbol{\theta}$ using $\nabla_{\boldsymbol{\theta}}\mathcal{L}$.
13:end for
14:// — Inference Phase —
[…]
17:Return $\mathbf{P}_{\text{final}}$
@@ 5.1 Setting
Datasets. We evaluate on standard TTA benchmarks [maharana2024texttt, osowiechi2024watt, hakim2024clipartt], covering two primary types of domain shift: i) Visual corruptions: CIFAR-10C, CIFAR-100C, ImageNet-C, and Tiny-ImageNet-C [datasets_corruptions]; and ii) Domain generalization: PACS [pacs], VLCS [vlcs], OfficeHome [officehome], and VisDA-C [visda]. We also use the original clean benchmarks (e.g., CIFAR-10 [cifar], CIFAR-10.1 [cifar101], and Tiny-ImageNet [TinyImagenet]) for ablation. Full dataset details are presented in Appendix A.
Baselines: We resort to relevant CLIP-based TTA parameter-efficient adaptation approaches, i.e., modify a set of learnable parameters, including TENT [wangtent], SAR [niutowards], VTE [dobler2024lost], TPT [shu2022test], CLIPArTT [hakim2024clipartt], WATT [osowiechi2024watt], and BATCLIP [maharana2024texttt], where [hakim2024clipartt, osowiechi2024watt, maharana2024texttt] represent the state-of-the-art.
Architectures: Following prior work, we use Vision Transformer (ViT) backbones from CLIP [radford2021learning]. Our main experiments use ViT-B/32 as in [hakim2024clipartt, osowiechi2024watt], and we demonstrate generalization with additional evaluations on larger ViT-B/16 and L/14 backbones, particularly CLIP’s and SigLIP’s [zhai2023sigmoidlosslanguageimage].
Implementation details: Following prior work [hakim2024clipartt, osowiechi2024watt], we use a batch size of 128 for all experiments (see Appendix B.4 for a sensitivity analysis on batch size). All reported results are the average of three runs using different random seeds. We follow the officially recommended hyperparameters for all baseline methods. Specifically, TENT [wangtent] and CLIPArTT [hakim2024clipartt] are run for 10 adaptation steps per batch. For WATT [osowiechi2024watt], we use their suggested $L=2$, $M=5$, and 8 templates, totaling 80 iterations per batch. For BATCLIP [maharana2024texttt], we adapt both the text and visual encoders. For all other baselines, we use the learning rates specified in their respective papers for each dataset. We set the learning rate for our method to $10^{-4}$ for all datasets, except for ImageNet-C, where we use $5\times 10^{-5}$ as in [maharana2024texttt]. Our formulation introduces only two new hyperparameters: the entropic constraint weight, $\epsilon$ from Eq. 7, and the number of Sinkhorn iterations, $T$. These are fixed across all datasets and experiments. We set $\epsilon=0.7$ (see Appendix B.3) and $T=3$. We found $T=3$ to be sufficient for convergence, as in prior work [caron2018deep].